OpenStudy (anonymous):

Solve for x? 2^x+6 = 5^x-2

OpenStudy (anonymous):

Just to clarify, the + 6 and -2 are not in the exponent, right? because if they are you need parentheses

OpenStudy (anonymous):

Oh yes. Sorry. They are in the exponent. I'll remember that for future reference. Thank you so much.

OpenStudy (anonymous):

It's been a while for me on logs so give me a sec

OpenStudy (anonymous):

Oh sure. That's quite fine. I just appreciate the help. :)

OpenStudy (lgbasallote):

so the question is \[\huge 2^{x+6} = 5^{x-2}\]right?

OpenStudy (anonymous):

I'll give this one to Igb

OpenStudy (lgbasallote):

nahh i just wanted to do some latex

OpenStudy (anonymous):

Yep, that's right. :) Thanks.

OpenStudy (anonymous):

So how would I go about solving it?

OpenStudy (lgbasallote):

\[\Large 2^{x+6} = 5^{x-2}\] take the log of both sides \[\Large \log 2^{x+6} = \log 5^{x-2}\] use power rule \[\Large(x+6) \log 2 = (x-2) \log 5\] isolate x \[\Large \frac{x+6}{x-2} = \frac{\log 5}{\log 2}\] use change of base formula \[\Large \frac{x+6}{x-2} = \log_2 5\] now solve log_2 5

OpenStudy (anonymous):

Thanks for the help. So how would I solve the next portion of this? I really don't know how to do this one at all. :(

OpenStudy (lgbasallote):

you. solve3. log_2 5. like. i. said.

OpenStudy (lgbasallote):

nevermind that 3 there. it's a typo

OpenStudy (anonymous):

Alright. So would it be 1.505149978?

OpenStudy (lgbasallote):

hmm no...

OpenStudy (lgbasallote):

what did your calculator tell you about \[\log 5 \div \log 2\]

OpenStudy (anonymous):

I tried putting that into my calculator now. Is it 2.321928095?

OpenStudy (lgbasallote):

yes. so \[\frac{x+6}{x-2} \approx 2\] \[\implies x + 6 \approx 2(x-2)\] \[\implies x + 6\approx 2x - 4\] \[\implies 4 + 6 \approx 2x - x\] \[\implies 10 \approx x\] do you follow that?

OpenStudy (anonymous):

I think so. So, the answer is 10 then? Or is there more to this problem?

OpenStudy (lgbasallote):

as far as im concerned.. i'd like it to be 10 period. although..it is a little inaccurate since you did logarithms but heh

OpenStudy (anonymous):

my answer comes out to be 0.398

OpenStudy (anonymous):

Thank you for your help too. Is that the answer you got too Igbasallote?